Deter the current that will be drawn from a 6v source connected to a 60mm length of wire, with a resistvity of0. 0185micro ohm meter. With area of wire given as 0.001meter square
1
Expert's answer
2019-12-06T09:40:03-0500

The resistance:


"R=\rho\frac{l}{A}"

The current:


"I=\frac{U}{R}=\frac{AU}{\rho l}"

"I=\frac{(0.001)6}{(0.0185\cdot 10^{-6})(0.06)}=5.4\cdot 10^{6}\ A"